OpenStudy (anonymous):

Can anyone help me please? a student is taking a test in which items of type A are worth 10 points and items of type B are worth 15 points. It takes 3 min to complete each item of type A and 8 min to complete each item of type B. The total time allowed is 60 min and Anna answers exactly 15 questions. How many questions of each type did she complete? Assuming that all her answers were correct, what was her score?

OpenStudy (anonymous):

A: # type A items B: # type B items "It takes 3 min to complete each item of type A and 8 min to complete each item of type B. The total time allowed is 60 min" ==>3A + 8B = 60 "Anna answers exactly 15 questions." ==> A + B = 15 You have a system of two equations. You can solve it by substitution. "items of type A are worth 10 points and items of type B are worth 15 points" ==> 10A + 15B = ? <== Use this to find the total score once you have found A, B from the system of equations.

OpenStudy (anonymous):

Thank you for your help! Have a good night.